If f from R to R is continuous and open, show that f is strictly monotone.
1
Expert's answer
2011-05-06T05:55:53-0400
If f is not monotone, then there are two distinct points a<b such that f(a)=f(b). Then by Weierstrass' theorem f has a maximum C and minimum c on the closed interval [a,b]. We can assume that at least one of the numbers f(C) or f(c), say f(C) differs from f(a)=f(b).

Then the image open interval
f( (a,b) ) = (f(a), C]
is not open, which contradicts to openness of f.